# Uncertainty relations for overcomplete measurements from generalized equiangular tight frames

### v1

@@ IV Conclusions
We have derived uncertainty relations for measurements assigned to a generalized equiangular measurement. Each of such POVMs is built from generalized equiangular tight frames that are complementary to one another. This concept recently proposed in [28] allows us to manipulate with the involved coefficients in several ways. For instance, there are generalized equiangular measurement for which three out of four defining parameters becomes independent of the measurement operator indices. The role of measurements with a special inner structure is well known in quantum information theory. Complete sets of MUBs and SIC-POVMs are probably most known in this regard. At the same time, the question of existence for such measurements seems to be very hard. To mitigate these questions, one may to weak some of the imposed conditions. Mutually unbiased measurements [26] and general SIC-POVMs [26] are examples of such a kind. The paper [28] proposes further development in this direction.
@@
The presented results follow from an estimation of the index of coincidence for POVMs of interest. In addition, several properties of information characteristics were used to derive uncertainty relations. The used approach is based on inequalities obtained by means of information diagrams. In more detail, these questions are considered in [54, 55]. Namely, generalized entropies of both the Rényi and Tsallis types were utilized. Uncertainty relations with probabilities of separate outcomes have also been addressed. Dealing with average characteristics involves generally unequal weights. The latter is a flip side of existing freedom to manipulate the defining parameters. All the new uncertainty relations are briefly exemplified. Like MUBs and SICs, overcomplete measurements from generalized equiangular tight frames can allow one to detect entanglement and steerability. Applications of the presented uncertainty relations for such purposes will be considered in future investigations.

### v3

@@ IV Conclusions
We have derived uncertainty relations for measurements assigned to a generalized equiangular measurement. Each of such POVMs is built from generalized equiangular tight frames that are complementary to one another. This concept recently proposed in [29] allows us to manipulate with the involved coefficients in several ways. For instance, there are generalized equiangular measurement for which three out of four defining parameters becomes independent of the measurement operator indices. The role of measurements with a special inner structure is well known in quantum information theory. Complete sets of MUBs and SIC-POVMs are probably most known in this regard. At the same time, the question of existence for such measurements seems to be very hard. To mitigate these questions, one can weak some of the imposed conditions. Mutually unbiased measurements [27] and general SIC-POVMs [27] are examples of such a kind. The paper [29] proposes further development in this direction.
@@
The presented results follow from an estimation of the index of coincidence for POVMs of interest. In addition, several properties of information characteristics were used to derive uncertainty relations. The used approach is based on inequalities obtained by means of information diagrams. In more detail, these questions are considered in [55, 56]. Namely, generalized entropies of both the Rényi and Tsallis types were utilized. Uncertainty relations with probabilities of separate outcomes have also been addressed. Dealing with average characteristics involves generally unequal weights. The latter is a flip side of existing freedom to manipulate the defining parameters. All the new uncertainty relations are briefly exemplified. Like MUBs and SICs, quantum measurements from generalized equiangular tight frames can allow one to detect entanglement and steerability. Applications of the presented uncertainty relations for such purposes will be considered in future investigations.
Appendix A A statement on indices of coincidence
In principle, the required technique was developed by the author of [29]. The key distinction of this appendix is that the considered sets of operators are assumed to be incomplete in general. Instead of the equality presented in [29], the following inequality takes place.
Lemma 1
Let ${\mathbb{E}}=\bigl{\{}{\mathcal{E}}_{\mu}\bigr{\}}$ be a set of $M$ generalized symmetric POVMs assigned to a generalized equiangular measurement due to (14), with the $\mu$-th index of coincidence
$I({\mathcal{E}}_{\mu};\boldsymbol{\rho})=\sum_{j=1}^{N_{\mu}}\mathrm{tr}({\mathsf{E}}_{\mu,j}\boldsymbol{\rho})^{2}\,.$ (84)
It holds that
$\sum_{\mu=1}^{M}\omega_{\mu}\,I({\mathcal{E}}_{\mu};\boldsymbol{\rho})\leq\frac{d\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}w_{\mu}\omega_{\mu}\,,$ (85)
where the numbers $\omega_{\mu}$ and $\mathcal{C}$ are defined by (34) and (35), respectively.
Proof. Instead of (85), we aim to prove the equivalent inequality
$\sum_{\mu=1}^{M}\frac{1}{a_{\mu}^{2}(b_{\mu}-c_{\mu})}\,\biggl{(}\,\sum_{j=1}^{N_{\mu}}\mathrm{tr}({\mathsf{P}}_{\mu,j}\boldsymbol{\rho})^{2}-\frac{a_{\mu}\gamma_{\mu}}{d}\biggr{)}\leq\mathrm{tr}(\boldsymbol{\rho}^{2})-f\ .$ (86)
Their equivalence can be checked immediately due to (14) and (15). Let us represent an arbitrary $\boldsymbol{\rho}$ as
$\boldsymbol{\rho}=\sum_{\mu=1}^{M}\sum_{j=1}^{N_{\mu}}\mathrm{tr}({\mathsf{P}}_{\mu,j}\boldsymbol{\rho}){\mathsf{G}}_{\mu,j}+\boldsymbol{\varpi}\,.$ (87)
For the given elements ${\mathsf{P}}_{\mu,j}$, the operators ${\mathsf{G}}_{\mu,j}$ read as [29]
[…]
By $\boldsymbol{\varpi}$, we mean in (87) potentially nonzero part of $\boldsymbol{\rho}$ such that $\mathrm{tr}({\mathsf{G}}_{\mu,j}\boldsymbol{\varpi})=0$ for all $\mu$ and $j$. Being depending on $\boldsymbol{\rho}$, this term does not appear in (86). To check (87), one writes
$\mathrm{tr}({\mathsf{P}}_{\mu,i}\boldsymbol{\rho})=\sum_{\nu=1}^{M}\sum_{j=1}^{N_{\nu}}\mathrm{tr}({\mathsf{P}}_{\nu,j}\boldsymbol{\rho})\,\mathrm{tr}({\mathsf{P}}_{\mu,i}{\mathsf{G}}_{\nu,j})\,.$ (89)
It needs to show that the right-hand side of this formula indeed reduces to its left-hand side. Doing some calculations finally results in the formulas
$\sum_{\nu\neq\mu}\sum_{j=1}^{N_{\nu}}\mathrm{tr}({\mathsf{P}}_{\nu,j}\boldsymbol{\rho})\,\mathrm{tr}({\mathsf{P}}_{\mu,i}{\mathsf{G}}_{\nu,j})$ $=\frac{(M-1)a_{\mu}}{Md}\ ,$ (90)
$\sum_{j=1}^{N_{\mu}}\mathrm{tr}({\mathsf{P}}_{\mu,j}\boldsymbol{\rho})\,\mathrm{tr}({\mathsf{P}}_{\mu,i}{\mathsf{G}}_{\mu,j})$ $=\mathrm{tr}({\mathsf{P}}_{\mu,i}\boldsymbol{\rho})-\frac{(M-1)a_{\mu}}{Md}\ .$ (91)
Combining (90) with (91) proves (89).
Due to definition of $\boldsymbol{\varpi}$, we see from (87) that
$\mathrm{tr}(\boldsymbol{\rho}^{2})\geq\sum_{\mu=1}^{M}\sum_{\nu=1}^{M}\sum_{i=1}^{N_{\mu}}\sum_{j=1}^{N_{\nu}}\mathrm{tr}({\mathsf{P}}_{\mu,i}\boldsymbol{\rho})\,\mathrm{tr}({\mathsf{P}}_{\nu,j}\boldsymbol{\rho})\,\mathrm{tr}\bigl{(}{\mathsf{G}}_{\mu,i}{\mathsf{G}}_{\nu,j}\bigr{)}\,.$ (92)
It also follows from (88) that
$\mathrm{tr}\bigl{(}{\mathsf{G}}_{\mu,i}{\mathsf{G}}_{\nu,j}\bigr{)}=\frac{\delta_{\mu\nu}\delta_{ij}a_{\mu}^{2}(b_{\mu}-c_{\mu})+\delta_{\mu\nu}a_{\mu}^{2}(c_{\mu}-f)}{a_{\mu}^{2}(b_{\mu}-c_{\mu})a_{\nu}^{2}(b_{\nu}-c_{\nu})}+\frac{f}{M^{2}\gamma_{\mu}\gamma_{\nu}}\ .$ (93)
Substituting (93) into the right-hand side of (92) leads to the different sums. By calculations, the sum with the last term of (93) reads as
$\frac{f}{M^{2}}\sum_{\mu=1}^{M}\,\frac{1}{\gamma_{\mu}}\sum_{i=1}^{N_{\mu}}\mathrm{tr}({\mathsf{P}}_{\mu,i}\boldsymbol{\rho})\sum_{\nu=1}^{M}\,\frac{1}{\gamma_{\nu}}\sum_{j=1}^{N_{\nu}}\mathrm{tr}({\mathsf{P}}_{\nu,j}\boldsymbol{\rho})=\frac{f}{M^{2}}\sum_{\mu=1}^{M}\sum_{\nu=1}^{M}1=f\,.$ (94)
Substituting the first fraction in the right-hand side of (93) into (92) allows one to rewrite this inequality as
$\mathrm{tr}(\boldsymbol{\rho}^{2})-f\geq\sum_{\mu=1}^{M}\sum_{j=1}^{N_{\mu}}\frac{\mathrm{tr}({\mathsf{P}}_{\mu,j}\boldsymbol{\rho})^{2}}{a_{\mu}^{2}(b_{\mu}-c_{\mu})}+\sum_{\mu=1}^{M}\frac{\gamma_{\mu}^{2}(c_{\mu}-f)}{a_{\mu}^{2}(b_{\mu}-c_{\mu})^{2}}=\sum_{\mu=1}^{M}\sum_{j=1}^{N_{\mu}}\frac{\mathrm{tr}({\mathsf{P}}_{\mu,j}\boldsymbol{\rho})^{2}}{a_{\mu}^{2}(b_{\mu}-c_{\mu})}-\frac{1}{d}\,\sum_{\mu=1}^{M}\frac{\gamma_{\mu}}{a_{\mu}(b_{\mu}-c_{\mu})}\ .$ (95)
Here, we omitted some details based on the use of (8). It is clear that (95) is equivalent to the aim (86). $\blacksquare$
The inequality (85) allows us to estimate the index of coincidence averaged with the weights (34). In special cases, it reduces to the results obtained previously, including a set of MUBs [70], MUMs [51] and $(N,M)$-POVMs [23]. In all these cases, the weights (34) turned out to be equal. We refrain from presenting the details here.